# KiDS-SBI: Simulation-Based Inference Analysis of KiDS-1000 Cosmic Shear

### v1

@@ 3.5 Intrinsic alignments: non-linear alignment model
To model this effect, we follow the methodology set out in Tessore et al. (2023). In this prescription, for a given intrinsic alignment model, we define an associated effective convergence field, $\kappa^{(i)}_{\mathrm{IA}}$. This $\kappa^{(i)}_{\mathrm{IA}}$ is a useful construct which describes the contribution to the observed weak lensing signal from IA under the Born approximation. Alternatively, an IA model could also directly be implemented into the dark matter distribution of the matter fields which could be an interesting avenue for future work.
@@
For simplicity and consistency with typical modelling assumptions in weak lensing surveys such as KiDS-1000 (Joachimi et al., 2021), we choose to model $\kappa^{(i)}_{\mathrm{IA}}$ with the Non-Linear Alignment (NLA) model (Catelan et al., 2001; Hirata et al., 2004; Bridle & King, 2007), so we consider tidal alignments and do not consider tidal torquing. The NLA model assumes that the bias in the shear signal from IAs is linearly dependent on the projected local tidal field. The ”non-linear” part of the NLA model then simply refers to the fact that it has been found that the modelling of IAs is more accurate when modelling the underlying large-scale structure using a non-linear matter power spectrum, rather than a linear one (Bridle & King 2007, similar to our approach described in Sect. 3.1). This means that it is also proportional to the local matter overdensity field, which is given by (Hirata et al., 2004),
@@
where $A_{\mathrm{IA}}$ is the intrinsic alignments amplitude (which we are treating as a nuisance parameter that is also sampled, see Sect. 5), $C_{1}$ is a normalisation constant which we set to $C_{1}=5\times 10^{-14}h^{-2}M_{\odot}^{-1}\mathrm{Mpc}^{3}$ in agreement with the IA measurements at low redshifts by SuperCOSMOS (Brown et al., 2002), $\overline{\rho}_{\mathrm{cr}}(\overline{z}^{(i)};\mathbf{\Theta})$ is the mean critical matter density as a function of redshift and $D(\overline{z}^{(i)};\mathbf{\Theta})$ is the linear growth factor normalised to be unity at $z=0$. Note that in Eq. (17), we do not consider an explicit redshift dependence which is often expressed as a power-law term (Joachimi et al., 2011). We omit this term, as it is not considered in the fiducial KiDS-1000 analysis (Asgari et al., 2021; van den Busch et al., 2022), and it has been found that at least for $z<1$, the IA signals do not vary largely with redshift (Fortuna et al., 2021).
Nevertheless, the NLA model still comes with some caveats. Mainly, it neglects non-linearities which may not be captured in the underlying matter power spectrum, such as source density weighing, i.e. the next-to-leading order perturbations proportional to the source density considered in a full tidal alignment model (Blazek et al., 2015). This effect can in fact produce a signal in the shape measurements comparable to the correction due to non-linear structure growth (Krause et al., 2016). Despite these limitations, the NLA model has been found to be accurate enough to model the IA signal seen by past and current weak lensing surveys (Joachimi et al., 2011; Blazek et al., 2011; Heymans et al., 2013; Krause et al., 2016; Hilbert et al., 2017), including KiDS-1000 (Fortuna et al., 2021).
@@ 3.6 Galaxy positions and redshifts
With the cosmological dependence of the underlying large-scale structure characterised and astrophysical effects such as intrinsic alignments and baryonic feedback included, the forward simulations still need to include observational biases from our instruments. We choose to apply these on the level of galaxies, i.e. by biasing the sampled galaxy positions, redshifts and shapes/shears in accordance with our models for the observational systematic effects.
When sampling the galaxies’ positions on the sky, we consider the survey footprint, $\Omega_{\mathrm{survey}}$, the galaxy number density which the survey can observe, $n_{\mathrm{gal}}$, and its variability as a function of redshift and observational depth across the sky, i.e. as a function of the local observing conditions. When sampling the galaxies’ redshifts, we take into consideration the redshift distributions calibrated from photometry for a given survey, $n(z)$. Lastly, when determining the observed galaxy shapes, we consider the intrinsic ellipticity dispersion of the galaxies, $\sigma_{\mathrm{\epsilon}}$, the multiplicative and additive shear bias, while also including the effects on the measured galaxy shapes caused by the variation in the point spread function of the instrument as a function of the position in the sky.
Since we are interested in simulating a photometric survey, we must model the resolution along the line-of-sight realistically. This means that rather than modelling large-scale structure observables across $N_{\mathrm{shells}}$ concentric shells, they ought to be modelled across $N_{\mathrm{tomo}}$ tomographic bins. The tomographic bins of a given survey are defined by finite usually non-overlapping domains in photometric redshift, $z_{\mathrm{ph}}$, in which all observed galaxies are grouped, such that $z_{p}<z_{\mathrm{ph}}\leq z_{p+1}$, $p\in\{1,2,3,...,N_{\mathrm{tomo}}\}$ [^12] Note that we use the indices $p$ and $q$ for tomographic bins and the indices $i$ and $j$ for underlying shells.. The measured photometric redshift of a galaxy does not necessarily agree with its true/spectroscopic redshift, $z$. This means that galaxies which are grouped within a given photometric redshift bin, $p$, may not be located in this redshift range. To account for this, one can characterise a conditional redshift distribution for each tomographic bin $P(z|z_{\mathrm{ph}})$ by mapping the measured $z_{\mathrm{ph}}$ to known spectroscopic redshift measurements which have a comparatively low bias in the redshift estimate (Hildebrandt et al., 2021; van den Busch et al., 2022). Therefore, a given galaxy is in an associated pixel, $m$, shell, $i$, and tomographic bin, $p$. The probability of the galaxy being in tomographic bin, $p$, while it is sampled in shell, $i$, then depends on $P(z|z_{\mathrm{ph}})$.
@@
where $|\Omega_{\mathrm{survey}}|$ is the effective area covered by the survey footprint, $\overline{M}^{(p)}$ is the mean multiplicative shear bias in tomographic bin $p$ as calibrated in (Giblin et al., 2021), and $w_{i}$ is the lensfit weight for a given galaxy $i$. Note that the effective galaxy density within the KiDS-1000 gold sample is not the same as the galaxy density, $n_{\mathrm{gal}}$, when not correcting for the lensfit weights for each galaxy. However, as KiDS-SBI does not model galaxy weights, in the forward simulations $n_{\mathrm{eff}}=n_{\mathrm{gal}}$ since $w_{i}=1\forall i$.
@@
where $\epsilon_{\mathrm{obs},1}$ and $\mathrm{i}\epsilon_{\mathrm{obs},2}$ are the two components of the observed galaxy ellipticity. Applying Eqs. (18) and (19) to the KiDS-1000 gold sample, we obtain the values shown in Table 1.
@@
Table 1: Parameters used to sample galaxies and their shapes in line with the expectations for the KiDS-1000 gold sample for each tomographic bin (from S1 to S5). $\overline{n}_{\mathrm{eff}}$ is the mean galaxy number density for a given tomographic bin, $\overline{\sigma}_{\epsilon}/\sqrt{2}$ is the mean per-component shape dispersion, $a_{n_{\mathrm{eff}}}$ and $b_{n_{\mathrm{eff}}}$ are the slope and y-intercept, respectively, for the linear interpolation of the galaxy density as a function of the root-mean-square of the background noise in the KiDS catalogue, $\sigma_{\mathrm{rms}}$, according to Eq. (28), while $a_{\sigma_{\epsilon}}$ and $b_{\sigma_{\epsilon}}$ are the parameters to linearly interpolate $\sigma_{\epsilon}$ from $\sigma_{\mathrm{rms}}$ according to Eq. (29).
Although it does not have detectable effects on the cosmic shear signal, the KiDS-SBI forward simulations sample galaxy positions such that they trace the underlying matter fields, so spatially varying systematics may be accurately applied. Using the SALMO [^13] Speedy Acquisition for Lensing and Matter Observables; https://github.com/Linc-tw/salmo framework (Joachimi et al., 2021) for simulating galaxy catalogues, we Poisson-sample galaxies within each HEALPix pixel, $m$, on the sky using the following expectation for the galaxy counts
@@
where $w_{m}(\Omega_{\mathrm{survey}})$ is the weight for a given pixel $m$ which we take to be unity if a pixel is within the survey’s mask/footprint, $\Omega_{\mathrm{survey}}$, and zero if it is not, $b^{(i)}$ is galaxy bias parameter within a given shell which we assume to be $b^{(i)}=1\forall i$ for the purposes of this cosmic shear analysis, $P_{m}(p|i)$ is the probability of a galaxy in pixel $m$ being detected within tomographic bin $p$ while being within shell $i$ (i.e. a discretised version of $P_{m}(z|z_{\mathrm{ph}})$), $n_{\mathrm{gal},m}^{(p)}$ is the observed galaxy number per unit area for a given tomographic bin $p$ and a given pixel $m$ (here the pixel dependence accounts for the spatial variability of the detection limit of galaxies as a function of observing conditions, the number of exposures, etc.; see Sect. 3.8), and $A_{\mathrm{pix},m}$ is the area of pixel $m$. As $n_{\mathrm{gal},m}^{(p)}$ changes due to variations in observational depth, it is more or less likely to observe high-redshift galaxies. Hence, a change in depth from one pixel to the next can change the observed redshift distribution (see Sect. 3.8).
@@
To sample each galaxy’s redshift, we can use the fact that we know exactly how many galaxies are in each shell $i$ for a given tomographic bin $p$ and pixel $m$ from $\smash{N_{m}^{(i)(p)}(\mathbf{\Theta})}$. We can therefore randomly assign an index $i$ and $p$ to each sampled galaxy in accordance with that number. We can then randomly sample a specific $z_{\mathrm{spec}}$ for each galaxy by assuming that the galaxies are uniformly distributed within a given shell. Again, this means that weak lensing and galaxy clustering measurements from the simulated galaxies do not contain any information about the large-scale structure along the line-of-sight below the scales of shells.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 5: The distributions for each tomographic bin (S1 to S5) from which the shear bias parameters from KiDS-1000 (Giblin et al., 2021) shown in Eq. (27) are sampled. The first panel from the left shows the multiplicative shear bias, $M^{(p)}$. The second and third panel show the real and imaginary part of the additive shear bias, $c_{1}^{(p)}$ and $c_{2}^{(p)}$, respectively. The fourth and fifth panel show the real and imaginary part of the amplitude of the shear bias due to variations in the point-spread function, $\alpha_{1}^{(p)}$ and $\alpha_{2}^{(p)}$, respectively.
@@ 3.7 Galaxy shapes
In order to determine the galaxy shapes and shears, we begin by assigning intrinsic ellipticities. We define a given galaxy’s shape through a complex ellipticity, $\epsilon$, and any galaxy ellipticity in this work as follows
@@
where $\mathrm{i}\equiv\sqrt{-1}$. The intrinsic $\epsilon_{1}$ and $\epsilon_{2}$ of each galaxy are then sampled within SALMO as two independent normal random variates with zero mean and a variance equal to $\sigma_{\epsilon,m}^{(p)2}$. As with the galaxy density, $n_{\mathrm{gal},m}^{(p)}$, the shape dispersion, $\sigma_{\epsilon,m}^{(p)}$, can also vary with tomographic bin and as a function of the position on the sky due to the effects of the spatial variability of observational depth. See Sect. 3.8 for a more detailed discussion. The ellipticity of each source galaxy is then altered through weak gravitational lensing by the matter along the line-of-sight. The exact shape distortion is given by (Seitz & Schneider, 1996; Hu, 2000),
@@
where $\epsilon_{\mathrm{lensed}}(\mathbf{\Theta})$ is the lensed galaxy ellipticity, $\epsilon_{\mathrm{int}}$ is the intrinsic galaxy ellipticity and $g(\mathbf{\Theta})$ is the reduced shear ($g\in\mathbb{C}$) which is given by
@@
where $\gamma(\mathbf{\Theta})$ is the shear factor ($\gamma\in\mathbb{C}$). The use of Eq. (23) is an extension to the KiDS-1000 analysis which assumed the reduced shear approximation, i.e. $\epsilon_{\mathrm{lensed}}\approx\epsilon_{\mathrm{int}}+g$. In any case, this should cause a negligible change in the measured shear considering the sensitivity of KiDS-1000 (Joachimi et al., 2021). To calculate the shear field, it is useful to decompose the spin-0 convergence field into spherical harmonics as follows
@@
where $\kappa_{\ell m}^{(i)}(\mathbf{\Theta})$ are the harmonic coefficients of the convergence field of a given shell $i$. One can then use the following relation to define the shear field’s harmonic coefficients, $\gamma_{\ell m}^{(i)}$, given by (Peebles, 1973),
@@
However, this lensed ellipticity, $\epsilon_{\mathrm{lensed}}(\mathbf{\Theta})$, is not the observed shape measurement in a weak gravitational lensing survey such as KiDS-1000. Additionally, the shape measurement may be distorted by instrumental effects or shape modelling inaccuracies. Some relevant effects are selection biases, noise biases from a low signal-to-noise ratio, biases in the galaxy weights, artefacts, non-linear CCD responses, asymmetries in the point spread function, etc. (see e.g. Mandelbaum 2018 for a review). To account for this, it is common to parametrically estimate any residual systematics which may be affecting the shape measurements (Hildebrandt et al., 2017, 2021; Zuntz et al., 2018; Giblin et al., 2021). Within the forward simulations, we use the first-order parametric expansion of the observed galaxy shapes given by Heymans et al. (2006) as follows
@@
where $i$ is the index for a single galaxy within tomographic bin $p$ and pixel $m$, $M^{(p)}$ is the multiplicative shear bias as measured for tomographic bin $p$, $\alpha^{(p)}$ gives the fraction of the PSF ellipticity which remains in the shear estimator, $\epsilon_{\mathrm{PSF}}(m)$ is the local PSF measured within pixel $m$ (see Figure 6 for a map of $\epsilon_{\mathrm{PSF}}$ in KiDS-1000), $\beta^{(p)}$ gives the amplitude of the shear bias due to residuals which are not taken into account by the PSF model, $\delta\epsilon_{\mathrm{PSF}}$ represents the residuals in question, and $c^{(p)}$ is the additive shear bias within a given tomographic bin $p$. Note that spatially varying multiplicative shear biases, $M^{(p)}$, can also arise, but previous work has found that the mean bias across all positions is sufficient to accurately model the effect (Kitching et al., 2019). In accordance with previous findings (Giblin et al., 2021), we take $\beta^{(p)}=0\,\forall p$. The amplitudes of the remaining systematics as calibrated by Giblin et al. (2021) are shown in Fig. 5. It becomes apparent that these effects are relatively small which explains why only the multiplicative and additive bias is taken into consideration in the modelling on the fiducial KiDS-1000 analysis (Giblin et al., 2021; Joachimi et al., 2021; Asgari et al., 2021; Heymans et al., 2021). Additionally, the measured values of $M^{(p)}$, $\alpha^{(p)}$ and $c^{(p)}$ have associated uncertainties. To take this uncertainty into account and make sure it is represented in the effective likelihood, rather than using a fixed value, we randomly sample different values for $M^{(p)}$, $\alpha^{(p)}$ and $c^{(p)}$ for each realisation of the forward simulations from a normal distribution as shown in Fig. 5. Note that we do not explicitly marginalise over these parameters, but as they are sampled randomly from predetermined probability density distributions, so that their associated uncertainty is propagated into the effective likelihood. This is similar to how the seed for each log-normal random field, and the galaxy positions which populate the associated matter fields are varied from one forward simulation to the next, such that cosmic variance enters the effective likelihood as well.

### v2

@@ 3.5 Intrinsic alignments: Non-linear alignment model
To model this effect, we followed the methodology set out in Tessore et al. (2023). In this prescription, for a given intrinsic alignment model, we define an associated effective convergence field, $\kappa^{(i)}_{\mathrm{IA}}$. This $\kappa^{(i)}_{\mathrm{IA}}$ is a useful construct which describes the contribution to the observed weak lensing signal from IA under the Born approximation. Alternatively, an IA model could also directly be implemented into the dark matter distribution of the matter fields which could be an interesting avenue for future work.
@@
For simplicity and consistency with typical modelling assumptions in weak lensing surveys such as KiDS-1000 (Joachimi et al., 2021), we chose to model $\kappa^{(i)}_{\mathrm{IA}}$ with the Non-Linear Alignment (NLA) model (Catelan et al., 2001; Hirata et al., 2004; Bridle & King, 2007), so we consider tidal alignments and do not consider tidal torquing. The NLA model assumes that the bias in the shear signal from IAs is linearly dependent on the projected local tidal field. The ’non-linear’ part of the NLA model then simply refers to the fact that it has been found that the modelling of IAs is more accurate when modelling the underlying large-scale structure using a non-linear matter power spectrum, rather than a linear one (Bridle & King 2007, similar to our approach described in Sect. 3.1). This means that it is also proportional to the local matter overdensity field, which is given by (Hirata et al., 2004),
@@
where $A_{\mathrm{IA}}$ is the intrinsic alignments amplitude (which we are treating as a nuisance parameter that is also sampled, see Sect. 5), $C_{1}$ is a normalisation constant which we set to $C_{1}=5\times 10^{-14}h^{-2}M_{\odot}^{-1}\mathrm{Mpc}^{3}$ in agreement with the IA measurements at low redshifts by SuperCOSMOS (Brown et al., 2002), $\overline{\rho}_{\mathrm{cr}}(\overline{z}^{(i)};\mathbf{\Theta})$ is the mean critical matter density as a function of redshift and $D(\overline{z}^{(i)};\mathbf{\Theta})$ is the linear growth factor normalised to be unity at $z=0$. We note that in Eq. (17), we do not consider an explicit redshift dependence which is often expressed as a power-law term (Joachimi et al., 2011). We omit this term, as it is not considered in the fiducial KiDS-1000 analysis (Asgari et al., 2021; van den Busch et al., 2022), and it has been found that at least for $z<1$, the IA signals do not vary largely with redshift (Fortuna et al., 2021).
Nevertheless, the NLA model still comes with some caveats. Mainly, it neglects non-linearities which may not be captured in the underlying matter power spectrum, such as source density weighing, expressly the next-to-leading order perturbations proportional to the source density considered in a full tidal alignment model (Blazek et al., 2015). This effect can in fact produce a signal in the shape measurements comparable to the correction due to non-linear structure growth (Krause et al., 2016). Despite these limitations, the NLA model has been found to be accurate enough to model the IA signal seen by past and current weak lensing surveys (Joachimi et al., 2011; Blazek et al., 2011; Heymans et al., 2013; Krause et al., 2016; Hilbert et al., 2017), including KiDS-1000 (Fortuna et al., 2021).
@@ 3.6 Galaxy positions and redshifts
With the cosmological dependence of the underlying large-scale structure characterised and astrophysical effects such as intrinsic alignments and baryonic feedback included, the forward simulations still need to include observational biases from our instruments. We chose to apply these on the level of galaxies, namely, by biasing the sampled galaxy positions, redshifts, and shapes and shears in accordance with our models for the observational systematic effects.
When sampling the galaxies’ positions on the sky, we consider the survey footprint, $\Omega_{\mathrm{survey}}$, the galaxy number density which the survey can observe, $n_{\mathrm{gal}}$, and its variability as a function of redshift and observational depth across the sky, namely, as a function of the local observing conditions. When sampling the galaxies’ redshifts, we take into consideration the redshift distributions calibrated from photometry for a given survey, $n(z)$. Lastly, when determining the observed galaxy shapes, we consider the intrinsic ellipticity dispersion of the galaxies, $\sigma_{\mathrm{\epsilon}}$, the multiplicative and additive shear bias, while also including the effects on the measured galaxy shapes caused by the variation in the point spread function of the instrument as a function of the position in the sky.
Since we are interested in simulating a photometric survey, we must model the resolution along the line of sight realistically. This means that rather than modelling large-scale structure observables across $N_{\mathrm{shells}}$ concentric shells, they ought to be modelled across $N_{\mathrm{tomo}}$ tomographic bins. The tomographic bins of a given survey are defined by finite usually non-overlapping domains in photometric redshift, $z_{\mathrm{ph}}$, in which all observed galaxies are grouped, such that $z_{p}<z_{\mathrm{ph}}\leq z_{p+1}$, $p\in\{1,2,3,...,N_{\mathrm{tomo}}\}$ [^12] Note: we use the indices $p$ and $q$ for tomographic bins and the indices $i$ and $j$ for underlying shells.. The measured photometric redshift of a galaxy does not necessarily agree with its true/spectroscopic redshift, $z$. This means that galaxies which are grouped within a given photometric redshift bin, $p$, may not be located in this redshift range. To account for this, we can characterise a conditional redshift distribution for each tomographic bin $P(z|z_{\mathrm{ph}})$ by mapping the measured $z_{\mathrm{ph}}$ to known spectroscopic redshift measurements which have a comparatively low bias in the redshift estimate (Hildebrandt et al., 2021; van den Busch et al., 2022). Therefore, a given galaxy is in an associated pixel, $m$, shell, $i$, and tomographic bin, $p$. The probability of the galaxy being in tomographic bin, $p$, while it is sampled in shell, $i$, then depends on $P(z|z_{\mathrm{ph}})$.
@@
where $|\Omega_{\mathrm{survey}}|$ is the effective area covered by the survey footprint, and $w_{i}$ is the lensfit weight for a given galaxy, $i$. We note that the effective galaxy density within the KiDS-1000 gold sample is not the same as the galaxy density, $n_{\mathrm{gal}}$, when not correcting for the lensfit weights for each galaxy. However, as KiDS-SBI does not model galaxy weights, in the forward simulations $n_{\mathrm{eff}}=n_{\mathrm{gal}}$ since $w_{i}=1\forall i$.
@@
where $\epsilon_{\mathrm{obs},i,1}$ and $\epsilon_{\mathrm{obs},i,2}$ are the two components of the observed galaxy ellipticity, and $\overline{M}^{(p)}$ is the mean multiplicative shear bias in tomographic bin $p$ as calibrated in (Giblin et al., 2021). Applying Eqs. (18) and (19) to the KiDS-1000 gold sample, we obtain the values shown in Table 13. We note that due to the way $\sigma_{\epsilon}$ is defined in Eq. (19), it incorporates both the noise due to intrinsic galaxy shapes as well as due to image noise in the survey.
Table 1: Parameters used to sample galaxies and their shapes in line with the expectations for the KiDS-1000 gold sample for each tomographic bin (from S1 to S5).
@@
[…]
Although it does not have detectable effects on the cosmic shear signal, the KiDS-SBI forward simulations sample galaxy positions such that they trace the underlying matter fields, so spatially varying systematics may be accurately applied. Using the SALMO [^14] Speedy Acquisition for Lensing and Matter Observables: https://github.com/Linc-tw/salmo framework (Joachimi et al., 2021) for simulating galaxy catalogues, we Poisson-sample galaxies within each HEALPix pixel, $m$, on the sky using the following expectation for the galaxy counts
@@
where $w_{m}(\Omega_{\mathrm{survey}})$ is the weight for a given pixel $m$ which we take to be unity if a pixel is within the survey’s mask/footprint, $\Omega_{\mathrm{survey}}$, and zero if it is not, $b^{(i)}$ is galaxy bias parameter within a given shell which we assume to be $b^{(i)}=1\forall i$ for the purposes of this cosmic shear analysis, $P_{m}(p|i)$ is the probability of a galaxy in pixel $m$ being detected within tomographic bin $p$ while being within shell $i$ (in other words, a discretised version of $P_{m}(z|z_{\mathrm{ph}})$), $n_{\mathrm{gal},m}^{(p)}$ is the observed galaxy number per unit area for a given tomographic bin $p$ and a given pixel $m$ (here the pixel dependence accounts for the spatial variability of the detection limit of galaxies as a function of observing conditions, the number of exposures, etc.; see Sect. 3.8), and $A_{\mathrm{pix},m}$ is the area of pixel $m$. As $n_{\mathrm{gal},m}^{(p)}$ changes due to variations in observational depth, it is more or less likely to observe high-redshift galaxies. Hence, a change in depth from one pixel to the next can change the observed redshift distribution (see Sect. 3.8).
@@
To sample each galaxy’s redshift, we can use the fact that we know exactly how many galaxies are in each shell $i$ for a given tomographic bin $p$ and pixel $m$ from $\smash{N_{m}^{(i)(p)}(\mathbf{\Theta})}$. We can therefore randomly assign an index $i$ and $p$ to each sampled galaxy in accordance with that number. We can then randomly sample a specific $z_{\mathrm{spec}}$ for each galaxy by assuming that the galaxies are uniformly distributed within a given shell. Again, this means that weak lensing and galaxy clustering measurements from the simulated galaxies do not contain any information about the large-scale structure along the line of sight below the scales of shells.
Figure 5: Distributions for each tomographic bin (S1 to S5) from which the shear bias parameters from KiDS-1000 (Giblin et al., 2021) shown in Eq. (27) are sampled. The first panel from the left shows the multiplicative shear bias, $M^{(p)}$. The second and third panel show the real and imaginary part of the additive shear bias, $c_{1}^{(p)}$ and $c_{2}^{(p)}$, respectively. The fourth and fifth panel show the real and imaginary part of the amplitude of the shear bias due to variations in the point-spread function, $\alpha_{1}^{(p)}$ and $\alpha_{2}^{(p)}$, respectively.
@@ 3.7 Galaxy shapes
To determine the galaxy shapes and shears, we begin by assigning intrinsic ellipticities. We define a given galaxy’s shape through a complex ellipticity, $\epsilon$, and any galaxy ellipticity in this work as follows
@@
where $\mathrm{i}\equiv\sqrt{-1}$. The intrinsic $\epsilon_{1}$ and $\epsilon_{2}$ of each galaxy are then sampled within SALMO as two independent normal random variates with zero mean and a variance equal to $\sigma_{\epsilon,m}^{(p)2}$. As with the galaxy density, $n_{\mathrm{gal},m}^{(p)}$, the shape dispersion, $\sigma_{\epsilon,m}^{(p)}$, can also vary with tomographic bin and as a function of the position on the sky due to the effects of the spatial variability of observational depth. See Sect. 3.8 for a more detailed discussion. The ellipticity of each source galaxy is then altered through weak gravitational lensing by the matter along the line of sight. The exact shape distortion is given by (Seitz & Schneider, 1996; Hu, 2000),
@@
where $\epsilon_{\mathrm{lensed}}(\mathbf{\Theta})$ is the lensed galaxy ellipticity, $\epsilon_{\mathrm{int}}$ is the intrinsic galaxy ellipticity and $g(\mathbf{\Theta})$ is the reduced shear ($g\in\mathbb{C}$), given by
@@
where $\gamma(\mathbf{\Theta})$ is the shear factor ($\gamma\in\mathbb{C}$). The use of Eq. (23) is an extension to the KiDS-1000 analysis which assumed the reduced shear approximation, where $\epsilon_{\mathrm{lensed}}\approx\epsilon_{\mathrm{int}}+g$. In any case, this should cause a negligible change in the measured shear considering the sensitivity of KiDS-1000 (Joachimi et al., 2021). To calculate the shear field, it is useful to decompose the spin-0 convergence field into spherical harmonics as follows:
@@
where $\kappa_{\ell m}^{(i)}(\mathbf{\Theta})$ are the harmonic coefficients of the convergence field of a given shell $i$. We can then use the following relation to define the shear field’s harmonic coefficients, $\gamma_{\ell m}^{(i)}$, given by (Peebles, 1973),
@@
However, this lensed ellipticity, $\epsilon_{\mathrm{lensed}}(\mathbf{\Theta})$, is not the observed shape measurement in a weak gravitational lensing survey such as KiDS-1000. Additionally, the shape measurement may be distorted by instrumental effects or shape modelling inaccuracies. Some relevant effects are selection biases, noise biases from a low signal-to-noise ratio, biases in the galaxy weights, artefacts, non-linear CCD responses, asymmetries in the point spread function, etc. (see for example Mandelbaum 2018 for a review). To account for this, it is common to parametrically estimate any residual systematics which may be affecting the shape measurements (Hildebrandt et al., 2017, 2021; Zuntz et al., 2018; Giblin et al., 2021). Within the forward simulations, we use the first-order parametric expansion of the observed galaxy shapes given by Heymans et al. (2006) as follows
@@
where $i$ is the index for a single galaxy within tomographic bin $p$ and pixel $m$, $M^{(p)}$ is the multiplicative shear bias as measured for tomographic bin, $p$, we have $\alpha^{(p)}$ as the fraction of the PSF ellipticity, which remains in the shear estimator, $\epsilon_{\mathrm{PSF}}(m)$ is the local PSF measured within pixel $m$ (see Fig. 6 for a map of $\epsilon_{\mathrm{PSF}}$ in KiDS-1000), $\beta^{(p)}$ gives the amplitude of the shear bias due to residuals which are not taken into account by the PSF model, $\delta\epsilon_{\mathrm{PSF}}$ represents the residuals in question, and $c^{(p)}$ is the additive shear bias within a given tomographic bin, $p$. We note that spatially varying multiplicative shear biases, $M^{(p)}$, can also arise, but previous work has found that the mean bias across all positions is sufficient to accurately model the effect (Kitching et al., 2019). In accordance with previous findings (Giblin et al., 2021), we take $\beta^{(p)}=0\,\forall p$. The amplitudes of the remaining systematics as calibrated by Giblin et al. (2021) are shown in Fig. 5. It becomes apparent that these effects are relatively small which explains why only the multiplicative and additive bias is taken into consideration in the modelling on the fiducial KiDS-1000 analysis (Giblin et al., 2021; Joachimi et al., 2021; Asgari et al., 2021; Heymans et al., 2021). Additionally, the measured values of $M^{(p)}$, $\alpha^{(p)}$, and $c^{(p)}$ have associated uncertainties. To take this uncertainty into account and make sure it is represented in the effective likelihood, rather than using a fixed value, we randomly sample different values for $M^{(p)}$, $\alpha^{(p)}$ and $c^{(p)}$ for each realisation of the forward simulations from a normal distribution as shown in Fig. 5. We note that we do not explicitly marginalise over these parameters, but as they are sampled randomly from predetermined probability density distributions, so that their associated uncertainty is propagated into the effective likelihood. This is similar to how the seed for each log-normal random field, and the galaxy positions which populate the associated matter fields are varied from one forward simulation to the next, such that cosmic variance enters the effective likelihood as well.